Please show me how to solve this problem:
2t + 4s = 10
3t + 3d = 9
1t + 3d + 2s = 8

Answer by stanbon(75887)

2t + 0d + 4s = 10
3t + 3d + 0s = 9
1t + 3d + 2s = 8
-------------------
Using the Matrix function of a TI caluculator I get:
t = 2
d = 1
s = 3/2
=================
Cheers,
Stan H.